OpenStudy (anonymous):

Can someone please solve and illustrate how to solve the following quadratic equations: y^2-4y-8; x^2+3x-10=0; and y^2-y-6? Separately of course, with illustrative steps for each. Please and thanks a bunch!

OpenStudy (mrnood):

Do you mean: y^2-4y-8=0 x^2+3x-10=0 y^2-y-6 =0 If so = what methods do you know for solving? Factorisation, Complete the square and Quadratic Formula are most used...

OpenStudy (anonymous):

factoring, were supposed to solve by factoring

OpenStudy (anonymous):

and yes, those are the problems that I am referring to.

OpenStudy (mrnood):

OK - first of all - don't be confused by the use of y in these equations This is just a shorthand for any unknown, and hence you can substitute x if it makes you things look more familiar.

OpenStudy (mrnood):

not all quadratrics can be solved by factorising (not factoring) do you know how to find out whether it can?

OpenStudy (anonymous):

A response using Mathematica v9 Home Edition is attached.

OpenStudy (anonymous):

I got your attachment and still am very confused

OpenStudy (anonymous):

MrNood no I do not

OpenStudy (mrnood):

the attachment doesn't help you to solve for yourself anyway. Do you understand the standard form of the quadratic? ax^2+bx+c = 0 (note - as I said - you can exchange y for x if oyu like...)

OpenStudy (anonymous):

just in case I wasn't clear, these equations are each separate examples given to me by my classmates to practice. They are not one big problem, nor are they linked together in anyway other than how they are supposed to be solved.

OpenStudy (mrnood):

I understand. so - to factorise you need (x+m)(x+n)=0 OK?

OpenStudy (anonymous):

k

OpenStudy (mrnood):

when you expand that it becomes x^2 +mx +nx + mn = x^2 +x(m+n) +mn OK? - please don't say OK if oyu don't follow.....

OpenStudy (anonymous):

yes I believe I follow so far

OpenStudy (mrnood):

so we are looking foe the 2 numbers m and n that give the equation we need to factorise. Take ift from me for now - the first one doesn't factorise SO the second one x^2 +3x-10 = (x+m)(x+n) you cne see from the expansion that m+n = 3 mn=-10 can you find 2 numbers that work for these equations? (It's mainly trial and error..)

OpenStudy (anonymous):

numbers that would work for what? to use for the variables?

OpenStudy (mrnood):

From the expansion above, compared to your equation you know that m+n = 3 mn=-10 Find m and n so these are true

OpenStudy (anonymous):

5 + -2 for the first and 5 times -2 for the second

OpenStudy (mrnood):

Good so (x+5)(x-2) = x^2 +3x -10 OK?

OpenStudy (anonymous):

ok

OpenStudy (mrnood):

so going back to the original equation: x^2 +3x -10=0 we can change it to (x+5)(x-2) =0 OK?

OpenStudy (anonymous):

ok

OpenStudy (mrnood):

so the 2 answers are x+5=0 x-2 = 0 solve for the 2 separate values of x....

OpenStudy (anonymous):

-5 for the first one and 2 for the second one

OpenStudy (mrnood):

That is correct. I have to go - but For the third question: y^2-y-6 =0 (y+m)(y+n)=y^2-y-6 SO m+n = -1 mn = -6 Do the same as above for the solution...

OpenStudy (mrnood):

You're welcome, don't mention it

OpenStudy (anonymous):

ok I am trying to write this out and explain it in a post for my class, and have no idea how to show my work of working out the problem. I understand what was said, and I understand the solution, but how to write it all out and illustrate the steps to solve?

OpenStudy (mrnood):

First - NOT ALL QUADRATICS CAN BE FACTORISED. the general form is ax^2 + bx +c =0 We will deal for now with the situation where a = 1: If it can be factorised then by definition x^2 + bx +c= (x+m)(x+n)=0 so if you expand the brackets ax^2+bx+c = x^2+ x(m+n) +mn so you can see that m+n= b mn = c SO we need to find 2 numbers m & n such tha tey multiply to make c, and add to make b. The best way to do this is to look at the factors of c, and find a pair that add up to b When you have found m and n you can write the equation (x+m)(x+n)=0 so either (x+m) = 0 or (x+n) =0 These two solutions for x are the solution for the original quadratic